# Machine Learning vs. SED-Fitting: A Comparative Analysis of Accuracy in Stellar Mass Estimation

### v1

@@ I Introduction
Stellar mass ($M_{*}$) is a fundamental property of galaxies, governing their evolution and interactions with the cosmic environment (e.g., bell2001stellar; kauffmann2003dependence; peng2010mass). Accurate measurements of $M_{*}$ are critical for understanding star formation histories (SFHs), feedback processes, and the assembly of large-scale structures (e.g., conroy2009propagation; behroozi2013average; somerville2015physical). While dynamical mass estimates are available for nearby galaxies, their reliance on resolved kinematics makes them challenging for high-redshift objects due to limited angular resolution and faintness. Consequently, most high-redshift studies rely on photometric methods, primarily spectral energy distribution (SED)-fitting (e.g., boss2000possible; conroy2013modeling).
@@
Despite their widespread use, traditional SED-fitting tools, such as LePhare (Arnouts1999; Ilbert2006),a widely used and well-documented code, face several well-documented challenges, including inherent uncertainties arising from necessary assumptions about a galaxy’s composition (e.g., Acquaviva2015; Pacifici2015; Iyer2017), degeneracies between age, dust, and metallicity (e.g., conroy2009propagation), sensitivity to assumed SFH parametrizations (e.g., carnall2018inferring), and computational inefficiency when applied to large, deep photometric surveys requiring extensive parameter space exploration (e.g., Hemmati2019; asadi2025leveraging).
@@
Machine learning (ML) approaches offer promising alternatives by learning non-linear relationships directly from data, helping to break classical degeneracies such as age-dust-metallicity (e.g., baron2019machine), achieving orders-of-magnitude speed improvements, and leveraging information from all photometric bands simultaneously (e.g., Hemmati2019; Davidzon2022; asadi2025leveraging; asadi2025semi; Asadi_2025). However, a critical consideration for ML models is that if they are trained on simulated data or idealized stellar populations that suffer from the same inherent degeneracies and biases as traditional methods, the ML model will similarly inherit these limitations.
@@
asadi2025leveraging demonstrated that various machine learning algorithms—including Parametric t-SNE (Policar2021, hereafter Pt-SNE), Principal Component Analysis (mackiewicz1993principal), K-means (macqueen1967some), Random Forest (Breiman2001), and Self-Organizing Maps (Kohonen1981)—can achieve stellar mass estimates for COSMOS galaxies (Laigle2016) in the redshift range $0.8<z<1.2$ with accuracy comparable to LePhare (Arnouts1999; Ilbert2006) when trained on synthetic simple stellar populations (Bruzual2003, hereafter BC03), while offering significant computational advantages. Their study used LePhare as the benchmark for accuracy comparisons, establishing ML methods as viable alternatives for mass estimation.
@@
Building on this foundation, we present a controlled comparison of ML and SED-fitting methods using a COSMOS-like galaxy catalog derived from the Horizon-AGN hydrodynamical simulation (dubois2014dancing; laigle2019horizon). The simulated catalog provides true stellar masses as a benchmark, enabling direct validation. We analyze a sample of simulated COSMOS-like galaxies, applying both LePhare (a traditional SED-fitting tool) and Pt-SNE (as a representative ML-based approach), with the latter trained on noise-injected BC03. This approach isolates methodological differences from template systematics while enabling direct validation against ground-truth masses.
@@ II.1 Mock Galaxies
This study utilizes a mock galaxy catalog derived from the Horizon-AGN hydrodynamical simulation [^1] http://www.horizon-simulation.org/ (dubois2014dancing). The catalog, presented in (laigle2019horizon), contains 789,354 galaxies extracted from a $1\times 1$ deg2 lightcone using the AdaptaHOP halo finder (aubert2004origin) applied to the stellar particle distribution. Each stellar particle (with mass $\sim 2\times 10^{6}M_{\odot}$) is associated with a synthetic simple stellar population using the BC03 models, assuming a Chabrier initial mass function (IMF) (Chap2003) and interpolating between the metallicity values available in BC03.
@@
The Horizon-AGN virtual observatory reproduces the optical and near-infrared (NIR) photometry of COSMOS2015 galaxies (Laigle2016), including:
@@
  • 14 medium-band filters (Subaru/SuprimeCam; (taniguchi2007cosmic))
@@
For each filter, the signal-to-noise ratio (SNR) distribution replicates that of the ultra-deep stripes of COSMOS2015. The reference $3\sigma$ detection limits (measured in 3′′ apertures) are $K_{s}<24.7$ and $i^{+}<26.2$ (complete sensitivity depths for all filters are provided in Table 1 of (Laigle2016)).
@@ III.1 Algorithm
This study employs Pt-SNE [^2] https://opentsne.readthedocs.io/en/latest/parameters.html (Policar2021), an advanced extension of the t-SNE algorithm (van2008visualizing), to analyze high-dimensional galaxy data. This method learns a neural network that maps high-dimensional inputs to a low-dimensional embedding space, which then serves as a reduced feature space for predicting physical properties like stellar mass. The algorithm preserves the local neighborhood structure of the data, which is crucial for subsequent property inference.
@@ III.2 Training Data
To train the Pt-SNE, we utilized BC03 stellar population synthesis code to generate galaxy spectra with different physical properties. The SED of modeled galaxies is constructed by combining simple stellar populations with a range of stellar ages ($7.7<\log_{10}(\rm age/yr)<10.0$), according to an exponentially declining SFH. The e-folding timescale, $\tau$, is adopted to vary in the range of 0.1-10 Gyr. For each simple stellar population, a sub-solar metallicity ($0.4Z_{\odot}$) is adopted and stars are assumed to form following the Chabrier IMF (Chap2003). Furthermore, we added Calzetti extinction (Calzetti2000) with varying degrees of reddening (0<E(BV)<10<E(B-V)<1) to the synthetic SEDs.
@@ III.3 Pre-processing
  • First, for each color in the COSMOS-like sample, we trained a RandomForestRegressor model [^3] https://scikit-learn.org (e.g., see macqueen1967some; biau2016random) using the color values as features and their corresponding color errors as targets.
@@ IV SED-fitting Method
Stellar mass estimates for the galaxies were extracted from the Horizon-AGN catalog (laigle2019horizon). These values were derived by the catalog’s creators through a SED-fitting analysis using the LePhare code. The fitting methodology followed the same procedure as the COSMOS2015 catalog (Laigle2016).
The analysis utilized the BC03 template library model, adopting a Chabrier IMF (Chap2003). The models considered two different stellar metallicities ($Z_{\odot}$ and $0.4Z_{\odot}$) and two extinction laws: arnouts2013encoding and Calzetti2000. For the star formation histories (SFHs), two analytic parameterizations were employed: Exponentially declining: $\mathrm{SFR}(t)\propto e^{-t/\tau}$ and Delayed (for gas infall modeling): $\mathrm{SFR}(t)\propto te^{-t/\tau}/\tau$. where $\tau$ represents the e-folding time. The following time-steps are used: $\tau$ = 0.1, 0.3, 1, 3, 4, and 30 Gyr for the exponentially declining SFHs, and $\tau$ = 1 and 3 Gyr for the delayed SFHs. In the delayed SFH case, $\tau$ also corresponds to the galaxy age at peak SFR. These models initiate star formation at $t=0$ and do not reproduce multiple bursts. Stellar population ages are sampled up to the age of the Universe at each galaxy’s redshift, with up to 44 time steps at $z=0$.
@@ V.5 Computational Performance Comparison
The speedup factor grows with sample size, reaching approximately $3.2\times 10^{3}$ times faster than LePhare on average. This scaling relation is consistent with the computational performance of Pt-SNE reported in asadi2025leveraging. For the largest dataset (107 galaxies), Pt-SNE completes in $\sim$26 minutes while LePhare would require $\sim$139 days. This dramatic difference highlights Pt-SNE’s suitability for large-scale analyses where traditional methods become computationally prohibitive.
@@ VI.1 Accuracy Comparison: The Bias Discrepancy
  • LePhare operates on a discrete, sparse grid of parameters (ages, metallicities, $\tau$). The true physical parameters of a simulated galaxy form a continuous distribution that will almost never coincide exactly with a grid point. During $\chi^{2}$ minimization, the algorithm is forced to select the nearest available template in this discrete space. This process introduces a systematic discretization error or quantization bias. Crucially, this error is not random; it depends on the location of the true parameters relative to the fixed grid. If the true parameter distribution is not uniformly centered within grid cells across the parameter space, the average of many such fits will exhibit a net bias. This is a known source of systematic error in grid-based Bayesian inference methods (e.g., brammer2008eazy). In principle, this grid-induced bias can be reduced by employing a significantly finer parameter grid. However, this mitigation strategy comes at a steep computational cost, as the number of templates—and thus the fitting time—scales with the product of the number of discrete values for each parameter. For large-scale surveys with millions of galaxies, this often makes an exhaustively fine grid computationally prohibitive, presenting a fundamental trade-off between accuracy and practical feasibility.
  • Even with identical underlying models (BC03), the notorious age-dust-metallicity degeneracy creates extended, near-flat valleys in the multi-dimensional $\chi^{2}$ surface. Within these valleys, many different parameter combinations yield similarly good fits to the photometry. In such a high-dimensional, degenerate likelihood landscape, the $\chi^{2}$ minimization (or Bayesian marginalization) does not converge to a unique solution centered on the truth. Instead, it can settle into any point within the degenerate valley. The final selected parameters are highly sensitive to the specific topography of the likelihood surface, noise realizations, and prior assumptions. This process systematically favors certain regions of parameter space (e.g., younger, dustier, and less massive solutions) over others, leading to a consistent bias in the ensemble of recovered masses (e.g., conroy2009propagation; Pacifici2015).
@@
  • ML algorithms excel at capturing complex, non-linear relationships between galaxy photometry and physical properties. Unlike SED-fitting, which is constrained to a specific family of physical models within its template library, ML methods can learn more flexible mappings from the data without being bound by pre-defined astrophysical parameterizations (e.g., baron2019machine; kembhavi2022machine).
  • ML methods can break degeneracies by learning in high-dimensional photometric space. Where SED-fitting might struggle to distinguish between, for example, young dusty galaxies and old metal-rich systems based on similar broad-band filters (e.g., leja2017deriving; leja2019older), an ML model can identify a unique solution by recognizing subtle, correlated patterns across the entire set of photometric bands simultaneously. This high-dimensional mapping allows it to separate galaxy types that appear degenerate in lower-dimensional projections.
@@
At high masses, increasing the SNR does not significantly improve either method’s performance, indicating that observational noise is no longer the dominant source of uncertainty (see Figure 5). LePhare’s accuracy deteriorates markedly in this regime due to a fundamental mismatch between its simplified stellar population assumptions and the complex physics of massive galaxies in the HORIZON-AGN simulation. LePhare assumes exponentially declining SFHs, while the mock catalog contains massive galaxies with diverse SFHs shaped by mergers, AGN feedback, and rejuvenation events (dubois2014dancing). Additionally, the assumed restricted metallicity options in the mock catalog’s existing LePhare configuration (laigle2019horizon) fail to fully capture broader metallicity ranges characteristic of massive galaxies (kaviraj2017horizon). Subtle degeneracies among age, metallicity, and dust become more prominent in these evolved, older stellar populations, which LePhare’s predefined templates struggle to disentangle, leading to less precise mass estimates. While Pt-SNE shows better performance, it is not immune to these limitations—having been trained on simple BC03 synthetic templates. This means Pt-SNE’s current accuracy is also constrained by the realism and diversity of its own training data. Future work could explore the impact of training Pt-SNE on more physically diverse and complex SFHs and metallicity distributions.
@@ VI.2 Computational Efficiency and Survey Scalability
Regarding execution time, LePhare is significantly slower than Pt-SNE. This is because LePhare analyzes each galaxy in our mock sample independently, comparing its spectrum to a vast library of pre-computed model galaxies through an iterative fitting process. This individual analysis and extensive comparison make it computationally intensive. In contrast, Pt-SNE requires only an initial fast training phase on the BC03 synthetic templates; afterward, the entire mock sample can be rapidly mapped onto its manifold, enabling efficient stellar mass estimation. By avoiding individual galaxy analysis and iterative fitting, Pt-SNE offers a significantly faster approach, making it more suitable for large-scale applications where speed is crucial (e.g., Hemmati2019; asadi2025leveraging).
@@ VI.3 Machine Learning Limitations abd Scope
For clarity and reproducibility, this study focused exclusively on Pt-SNE as our ML benchmark, providing a controlled comparison against the established LePhare SED-fitting pipeline. While other ML approaches—including manifold learning algorithms (e.g., Self-Organizing Maps, UMAP (mcinnes2018umap)), regression methods (e.g., Random Forest, KNN (peterson2009k)), and clustering techniques (e.g., HDBSCAN (mcinnes2017hdbscan), K-means)—could achieve comparable performance with similar data-driven advantages, our single-algorithm framework ensures unambiguous interpretation of results.
A notable limitation of ML approaches like Pt-SNE is their reduced interpretability compared to traditional SED-fitting. While SED-fitting explicitly links observed photometry to physical parameters (e.g., star formation history, metallicity, dust attenuation) through well-defined stellar population models, ML methods operate as “black boxes”, learning complex mappings without immediately transparent physical explanations for individual parameter derivations (e.g., meher2021deep; sen2022astronomical).

### v2

# Machine Learning vs. Spectral Energy Distribution Fitting: A Comparative Analysis of Accuracy in Stellar Mass Estimation

@@ I Introduction
Stellar mass ($M_{*}$) is a fundamental property of galaxies, governing their evolution and interactions with the cosmic environment (e.g., Bell and de Jong, 2001; Kauffmann et al., 2003; Peng et al., 2010). Accurate measurements of $M_{*}$ are critical for understanding star formation histories (SFHs), feedback processes, and the assembly of large-scale structures (e.g., Conroy et al., 2009; Behroozi et al., 2013; Somerville and Davé, 2015). While dynamical mass estimates are available for nearby galaxies, their reliance on resolved kinematics makes them challenging for high-redshift objects due to limited angular resolution and faintness. Consequently, most high-redshift studies rely on photometric methods, primarily spectral energy distribution (SED)-fitting (e.g., Boss, 2000; Conroy, 2013).
@@
Despite their widespread use, traditional SED-fitting tools, such as LePhare (Arnouts et al., 1999; Ilbert et al., 2006),a widely used and well-documented code, face several well-documented challenges, including inherent uncertainties arising from necessary assumptions about a galaxy’s composition (e.g., Acquaviva et al., 2015; Pacifici et al., 2015; Iyer and Gawiser, 2017), degeneracies between age, dust, and metallicity (e.g., Conroy et al., 2009), sensitivity to assumed SFH parametrizations (e.g., Carnall et al., 2018), and computational inefficiency when applied to large, deep photometric surveys requiring extensive parameter space exploration (e.g., Hemmati et al., 2019; Asadi et al., 2025c).
@@
Machine learning (ML) approaches offer promising alternatives by learning non-linear relationships directly from data, helping to break classical degeneracies such as age-dust-metallicity (e.g., Baron, 2019), achieving orders-of-magnitude speed improvements, and leveraging information from all photometric bands simultaneously (e.g., Hemmati et al., 2019; Davidzon et al., 2022; Asadi et al., 2025c, b, a). However, a critical consideration for ML models is that if they are trained on simulated data or idealized stellar populations that suffer from the same inherent degeneracies and biases as traditional methods, the ML model will similarly inherit these limitations.
@@
Asadi et al. (2025c) demonstrated that various machine learning algorithms—including Parametric t-SNE (Policar et al., 2021, hereafter Pt-SNE), Principal Component Analysis (Maćkiewicz and Ratajczak, 1993), K-means (MacQueen, 1967), Random Forest (Breiman, 2001), and Self-Organizing Maps (Kohonen, 1981)—can achieve stellar mass estimates for COSMOS galaxies (Laigle et al., 2016) in the redshift range $0.8<z<1.2$ with accuracy comparable to LePhare (Arnouts et al., 1999; Ilbert et al., 2006) when trained on synthetic simple stellar populations (Bruzual and Charlot, 2003, hereafter BC03), while offering significant computational advantages. Their study used LePhare as the benchmark for accuracy comparisons, establishing ML methods as viable alternatives for mass estimation.
@@
Building on this foundation, we present a controlled comparison of ML and SED-fitting methods using a COSMOS-like galaxy catalog derived from the Horizon-AGN hydrodynamical simulation (Dubois et al., 2014; Laigle et al., 2019). The simulated catalog provides true stellar masses as a benchmark, enabling direct validation. We analyze a sample of simulated COSMOS-like galaxies, applying both LePhare (a traditional SED-fitting tool) and Pt-SNE (as a representative ML-based approach), with the latter trained on noise-injected BC03. This approach isolates methodological differences from template systematics while enabling direct validation against ground-truth masses.
@@ II.1 Mock Galaxies
This study utilizes a mock galaxy catalog derived from the Horizon-AGN hydrodynamical simulation [^1] http://www.horizon-simulation.org/ (Dubois et al., 2014). The catalog, presented in (Laigle et al., 2019), contains 789,354 galaxies extracted from a $1\times 1$ deg2 lightcone using the AdaptaHOP halo finder (Aubert et al., 2004) applied to the stellar particle distribution. Each stellar particle (with mass $\sim 2\times 10^{6}M_{\odot}$) is associated with a synthetic simple stellar population using the BC03 models, assuming a Chabrier initial mass function (IMF) (Chabrier, 2003) and interpolating between the metallicity values available in BC03.
@@
The Horizon-AGN virtual observatory reproduces the optical and near-infrared (NIR) photometry of COSMOS2015 galaxies (Laigle et al., 2016), including:
@@
  • 14 medium-band filters (Subaru/SuprimeCam; (Taniguchi et al., 2007))
@@
For each filter, the signal-to-noise ratio (SNR) distribution replicates that of the ultra-deep stripes of COSMOS2015. The reference $3\sigma$ detection limits (measured in 3′′ apertures) are $K_{s}<24.7$ and $i^{+}<26.2$ (complete sensitivity depths for all filters are provided in Table 1 of (Laigle et al., 2016)).
@@ III.1 Algorithm
This study employs Pt-SNE [^2] https://opentsne.readthedocs.io/en/latest/parameters.html (Policar et al., 2021), an advanced extension of the t-SNE algorithm (der Maaten et al., 2008), to analyze high-dimensional galaxy data. This method learns a neural network that maps high-dimensional inputs to a low-dimensional embedding space, which then serves as a reduced feature space for predicting physical properties like stellar mass. The algorithm preserves the local neighborhood structure of the data, which is crucial for subsequent property inference.
@@ III.2 Training Data
To train the Pt-SNE, we utilized BC03 stellar population synthesis code to generate galaxy spectra with different physical properties. The SED of modeled galaxies is constructed by combining simple stellar populations with a range of stellar ages ($7.7<\log_{10}(\rm age/yr)<10.0$), according to an exponentially declining SFH. The e-folding timescale, $\tau$, is adopted to vary in the range of 0.1-10 Gyr. For each simple stellar population, a sub-solar metallicity ($0.4Z_{\odot}$) is adopted and stars are assumed to form following the Chabrier IMF (Chabrier, 2003). Furthermore, we added Calzetti extinction (Calzetti et al., 2000) with varying degrees of reddening (0<E(BV)<10<E(B-V)<1) to the synthetic SEDs.
@@ III.3 Pre-processing
  • First, for each color in the COSMOS-like sample, we trained a RandomForestRegressor model [^3] https://scikit-learn.org (e.g., see MacQueen, 1967; Biau and Scornet, 2016) using the color values as features and their corresponding color errors as targets.
@@ IV SED-fitting Method
Stellar mass estimates for the galaxies were extracted from the Horizon-AGN catalog (Laigle et al., 2019). These values were derived by the catalog’s creators through a SED-fitting analysis using the LePhare code. The fitting methodology followed the same procedure as the COSMOS2015 catalog (Laigle et al., 2016).
The analysis utilized the BC03 template library model, adopting a Chabrier IMF (Chabrier, 2003). The models considered two different stellar metallicities ($Z_{\odot}$ and $0.4Z_{\odot}$) and two extinction laws: Arnouts et al. (2013) and Calzetti et al. (2000). For the star formation histories (SFHs), two analytic parameterizations were employed: Exponentially declining: $\mathrm{SFR}(t)\propto e^{-t/\tau}$ and Delayed (for gas infall modeling): $\mathrm{SFR}(t)\propto te^{-t/\tau}/\tau$. where $\tau$ represents the e-folding time. The following time-steps are used: $\tau$ = 0.1, 0.3, 1, 3, 4, and 30 Gyr for the exponentially declining SFHs, and $\tau$ = 1 and 3 Gyr for the delayed SFHs. In the delayed SFH case, $\tau$ also corresponds to the galaxy age at peak SFR. These models initiate star formation at $t=0$ and do not reproduce multiple bursts. Stellar population ages are sampled up to the age of the Universe at each galaxy’s redshift, with up to 44 time steps at $z=0$.
@@ V.5 Computational Performance Comparison
The speedup factor grows with sample size, reaching approximately $3.2\times 10^{3}$ times faster than LePhare on average. This scaling relation is consistent with the computational performance of Pt-SNE reported in Asadi et al. (2025c). For the largest dataset (107 galaxies), Pt-SNE completes in $\sim$26 minutes while LePhare would require $\sim$139 days. This dramatic difference highlights Pt-SNE’s suitability for large-scale analyses where traditional methods become computationally prohibitive.
@@ VI.1 Accuracy Comparison: The Bias Discrepancy
  • LePhare operates on a discrete, sparse grid of parameters (ages, metallicities, $\tau$). The true physical parameters of a simulated galaxy form a continuous distribution that will almost never coincide exactly with a grid point. During $\chi^{2}$ minimization, the algorithm is forced to select the nearest available template in this discrete space. This process introduces a systematic discretization error or quantization bias. Crucially, this error is not random; it depends on the location of the true parameters relative to the fixed grid. If the true parameter distribution is not uniformly centered within grid cells across the parameter space, the average of many such fits will exhibit a net bias. This is a known source of systematic error in grid-based Bayesian inference methods (e.g., Brammer et al., 2008). In principle, this grid-induced bias can be reduced by employing a significantly finer parameter grid. However, this mitigation strategy comes at a steep computational cost, as the number of templates—and thus the fitting time—scales with the product of the number of discrete values for each parameter. For large-scale surveys with millions of galaxies, this often makes an exhaustively fine grid computationally prohibitive, presenting a fundamental trade-off between accuracy and practical feasibility.
  • Even with identical underlying models (BC03), the notorious age-dust-metallicity degeneracy creates extended, near-flat valleys in the multi-dimensional $\chi^{2}$ surface. Within these valleys, many different parameter combinations yield similarly good fits to the photometry. In such a high-dimensional, degenerate likelihood landscape, the $\chi^{2}$ minimization (or Bayesian marginalization) does not converge to a unique solution centered on the truth. Instead, it can settle into any point within the degenerate valley. The final selected parameters are highly sensitive to the specific topography of the likelihood surface, noise realizations, and prior assumptions. This process systematically favors certain regions of parameter space (e.g., younger, dustier, and less massive solutions) over others, leading to a consistent bias in the ensemble of recovered masses (e.g., Conroy et al., 2009; Pacifici et al., 2015).
@@
  • ML algorithms excel at capturing complex, non-linear relationships between galaxy photometry and physical properties. Unlike SED-fitting, which is constrained to a specific family of physical models within its template library, ML methods can learn more flexible mappings from the data without being bound by pre-defined astrophysical parameterizations (e.g., Baron, 2019; Kembhavi and Pattnaik, 2022).
  • ML methods can break degeneracies by learning in high-dimensional photometric space. Where SED-fitting might struggle to distinguish between, for example, young dusty galaxies and old metal-rich systems based on similar broad-band filters (e.g., Leja et al., 2017, 2019), an ML model can identify a unique solution by recognizing subtle, correlated patterns across the entire set of photometric bands simultaneously. This high-dimensional mapping allows it to separate galaxy types that appear degenerate in lower-dimensional projections.
@@
At high masses, increasing the SNR does not significantly improve either method’s performance, indicating that observational noise is no longer the dominant source of uncertainty (see Figure 5). LePhare’s accuracy deteriorates markedly in this regime due to a fundamental mismatch between its simplified stellar population assumptions and the complex physics of massive galaxies in the HORIZON-AGN simulation. LePhare assumes exponentially declining SFHs, while the mock catalog contains massive galaxies with diverse SFHs shaped by mergers, AGN feedback, and rejuvenation events (Dubois et al., 2014). Additionally, the assumed restricted metallicity options in the mock catalog’s existing LePhare configuration (Laigle et al., 2019) fail to fully capture broader metallicity ranges characteristic of massive galaxies (Kaviraj et al., 2017). Subtle degeneracies among age, metallicity, and dust become more prominent in these evolved, older stellar populations, which LePhare’s predefined templates struggle to disentangle, leading to less precise mass estimates. While Pt-SNE shows better performance, it is not immune to these limitations—having been trained on simple BC03 synthetic templates. This means Pt-SNE’s current accuracy is also constrained by the realism and diversity of its own training data. Future work could explore the impact of training Pt-SNE on more physically diverse and complex SFHs and metallicity distributions.
@@ VI.2 Computational Efficiency and Survey Scalability
Regarding execution time, LePhare is significantly slower than Pt-SNE. This is because LePhare analyzes each galaxy in our mock sample independently, comparing its spectrum to a vast library of pre-computed model galaxies through an iterative fitting process. This individual analysis and extensive comparison make it computationally intensive. In contrast, Pt-SNE requires only an initial fast training phase on the BC03 synthetic templates; afterward, the entire mock sample can be rapidly mapped onto its manifold, enabling efficient stellar mass estimation. By avoiding individual galaxy analysis and iterative fitting, Pt-SNE offers a significantly faster approach, making it more suitable for large-scale applications where speed is crucial (e.g., Hemmati et al., 2019; Asadi et al., 2025c).
@@ VI.3 Machine Learning Limitations abd Scope
For clarity and reproducibility, this study focused exclusively on Pt-SNE as our ML benchmark, providing a controlled comparison against the established LePhare SED-fitting pipeline. While other ML approaches—including manifold learning algorithms (e.g., Self-Organizing Maps, UMAP (McInnes et al., 2018)), regression methods (e.g., Random Forest, KNN (Peterson, 2009)), and clustering techniques (e.g., HDBSCAN (McInnes, 2017), K-means)—could achieve comparable performance with similar data-driven advantages, our single-algorithm framework ensures unambiguous interpretation of results.
A notable limitation of ML approaches like Pt-SNE is their reduced interpretability compared to traditional SED-fitting. While SED-fitting explicitly links observed photometry to physical parameters (e.g., star formation history, metallicity, dust attenuation) through well-defined stellar population models, ML methods operate as “black boxes”, learning complex mappings without immediately transparent physical explanations for individual parameter derivations (e.g., Meher and Panda, 2021; Sen et al., 2022).